checked status of cooling and power (with Micheal and Doug):
power was already connected and 110V-related stuff was already on
TSS was disabled via hardware switch: it has been kept off. TSS has to be switched on before opening first time the dome
HP UMAC was switched off via hardware switch: it has been kept off
Fans were switched off: fans have been switched on
cooling was connected but valves were closed:
when valves have been opened the inlet Swagelok connectors popped out, it was not correctly inserted. Swagelok reconnected properly, valves opened and checked no leakage at the Hub interface for few minutes.
coolant total flux: 20 liters/1.33min = 15 l/mn (mirror+cabinet). Flow in the mirror is 12.7 l/min (measured later after switching on the mirror unit). Measurements with the telescope horizon pointing.
ethernet and fastlink communication: OK
2 actuators of the outer ring have incorrect reading of capacitive sensors (2meter gap). Actuators are #260 and #261 (DSP ordering). The two actuators are right below one of the brackets of the mirror cover, possibly ribbon cables of the actuators have been moved during bracket installation. We (with Michael) used the scissor-lift to remove the bracket, and pushed the ribbon cables into the connectors through the shroud . Actuators restarted working properly and no misreading has been experienced even tapping on the cables to check signal stability.
no gap contamination
1h after powering on, housekeeping warnings for BcuStratix of crate 2 and 3 (45/46C vs 42C of the other crate-BCUs): no real problem, temperature in the chamber is high.
no possible testing the setting of the shell: telescope at horizon all day long